Electron in a certain energy level n=n ,can emit 3spectral line. When they are in another energy level n=n2, they can emit 6 spectral line,the orbital speed of electron in two orbit are in ratio

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
 
As
n(n-1)/2 = N ( no of spectral lines)
Hence
n1 = 3 
n2 = 4
And also we know that
V ?\LARGE \alpha 1/n
Hence
V1/ V2 = 4/3
Regards

Arun